OpenStudy (anonymous):

Given vector a = 2i + 2j + k and c = pi + qj , find value p and q for which c is a unit vector perpendicular to a

OpenStudy (anonymous):

do you know how to take the dot product of two vectors?

OpenStudy (anonymous):

Yup, but I need another equation

OpenStudy (anonymous):

c = pi + qj

OpenStudy (anonymous):

Does magnitude of c equal to one ?

OpenStudy (anonymous):

it should

OpenStudy (anonymous):

p^2 + q^2 = 1 ---> (c is unit vector) 2p+2q=0 => p=-q (as 'a' dot product 'c' = 0)